# A quantum information theoretic analysis of reinforcement learning-assisted quantum architecture search

### v2

@@ III.2.1 Contribution of individual qubits for change in entanglement
We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Figure 5. We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$. This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement. In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases. However, no such pattern is observed for the region of $\Delta_{q_{0}}^{q_{1}}<0$. When $\Delta_{c}>0.4$, the value of $\Delta_{q_{0}}^{q_{1}}$ monotonically decreases with increase in $\Delta_{c}$. Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
III.2.2 Problem with variational quantum state diagonalisation
@@
We present the main observations in Figure 6, where we plot the four eigenvalues of $21$, $2$-qubit Haar random mixed quantum states with the correlation in conditional entropy among the qubits. For an individual quantum state, the RL-agent proposes $\mathbf{s}$ different structures of admissible ansatzes. Then we get $\rho_{i}^{\prime}$ states where $i\in[0,\mathbf{s}]$. Afterwards, we calculate the conditional entropy of individual qubits for each state and take the median. [^1] Calculating the median over the mean is justified by the broad range of conditional entropy distribution, which makes the median a more appropriate measure of central tendency. Then, we calculate the correction in conditional entropy among the qubits using PCC.
@@
In Figure 6, we observe that a mild anti-correlation between the qubits is required to find the two largest eigenvalues. As the magnitude of the largest eigenvalues decreases, the mild anti-correlation turns into a strong correlation between the qubits. This indicates a strong correlation between the qubits is required to achieve the two largest eigenvalues. Meanwhile, we require a strong anti-correlation between the qubits to find the smallest eigenvalues of the same states. Then, as the magnitude of the smallest eigenvalue increases, the strong anti-correlation turns into mild anti-correlation.
@@
Figure 6: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the second largest eigenvalues increases. Meanwhile, the same correlation between the qubits increases with the magnitude of the third and the smallest eigenvalues. The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations. Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues. The correlation of conditional entropy between qubits is calculated using Pearson correlation coefficient (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent. The $\bigstar$ signifies the eigenvalues corresponding to the same state.
@@
We recall that the correlation is calculated among the qubits based on the conditional entropy of individual qubits. As it is not possible to have two different kinds (such as mild anti-correlation and strong anti-correlation at the same time) of correlation between the qubits in the same ansatz structure, we can not find the largest and smallest eigenvalues utilising the same ansatz. For example, we mark the eigenvalues of a state with $\bigstar$, where we observe that to find the largest two eigenvalues, the RL-ansatz need to induce a mild anti-correlation between the two qubits; meanwhile, to obtain the smallest eigenvalues, we require strong anti-correlation.
On the contrary, for the algorithm, as the agent adds a new gate to the RL-ansatz, the correlation in conditional entropy among the qubits increases. That is why it is easier for the RL-ansatz to propose admissible ansatz with mild anti-correlation. We further certify this observation in Figure 7, where we observe the correlation in conditional entropy between qubits for $21$ random quantum states. We see that most of the admissible ansatz provide a mild anti-correlation among the qubits, indicating that the VQSD algorithm primarily focuses on finding the largest eigenvalues rather than the smallest ones. Hence, it is easier for the variational diagonalisation algorithm to find the largest eigenvalues more efficiently than the smallest ones.
@@
Figure 7: The correlation in conditional entropy among qubits is in the range of mild correlation for $21$ random quantum states. This observation, along with the observation in Figure 6 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues. The average correlation among all the $21$ states is $-0.4263$ (horizontal black line), which is in the regime of mild anti-correlation.
IV Discussion
@@
We observe that the concurrence of the admissible ansatz ranges between an upper and lower bound, which are independent of the initialisation of the weights of the deep-Q network. The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement. The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase change point at concurrence $0.322$. We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound. These observations provide insight regarding the relation between the entanglement of the initial state and the entanglement generated by the RL-VQSD admissible ansatzes. We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem. Specifically, we propose an entanglement enhancing initial block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$ Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
@@
Furthermore, we quantify the contribution of each qubit in the admissible ansatz using conditional quantum entropy. For the admissible RL-VQSD ansatz, we observe an equal contribution from each qubit for the full range of change in the entanglement of the state due to the ansatz. Focusing on the task of obtaining the eigenvalues of the starting state, we observe that a strong correlation between the qubits of the admissible ansatz is required to obtain the two largest eigenvalues, while strong anti-correlation between the qubits is needed to find the smallest eigenvalues of the same state. Noting that it is impossible to have both strong correlation and anti-correlation between the qubits of the same ansatz, we conclude that we cannot simultaneously obtain both the largest and smallest eigenvalues from the RL-VQSD admissible ansatz. From the observation, most of the admissible ansatz has mild anti-correlation among the qubits, which explains why it is easier for the diagonalisation algorithm to find the largest eigenvalues than the smallest ones.
@@
For future work, it would be interesting to conduct more comprehensive study of the implication of introducing entanglement Horodecki et al. (2009) or nonlocality Brunner et al. (2014); Sadhu and Das (2023)-based reward functions of the RL-agent in QAS. Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations. Also, we may expect this to provide an optimal quantum circuit that provides both the lowest and highest eigenvalues of the input quantum state.
@@ A.2 Conditional quantum entropy
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, consider a purification |$\ket{\psi}_{ABE}\in\mathscr{H}_{A}\otimes\mathscr{H}_{B}\mathscr{H}_{E}$. It then follows that:
@@
Theorem 1.
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, the following inequality applies to the marginal entropy $S_{A}^{\rho^{\prime}_{AB}}$ and the conditional quantum entropy $S_{A|B}^{\rho^{\prime}_{AB}}$:
$S_{A}^{\rho^{\prime}_{AB}}\geq S_{A|B}^{\rho^{\prime}_{AB}}.$ (8)
The above theorem states that conditioning cannot increase entropy. The value of conditional entropy can be positive or negative depending on the quantum state for which it is evaluated. However, it cannot be arbitrarily large or arbitrarily small. The following theorem places an upper bound on the absolute value of the conditional quantum entropy.
Theorem 2.
The conditional entropy is bounded as follows:
$|S_{A|B}^{\rho^{\prime}_{AB}}|\leq\log\dim(\mathscr{H}_{A})$ (9)
The bound is saturated for $\rho^{\prime}_{AB}=\mathbbm{1}_{A}\otimes\sigma_{B}$, where $\mathbbm{1}_{A}$ is the maximally mixed state and $\sigma_{B}\in\mathscr{D}(\mathscr{H}_{B})$, and also for the family of maximally entangled states.
@@ A.3 Correlation coefficient
Appendix B The reinforcement learning enhanced variational quantum state diagonalisation
where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles. Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilised in the training process. (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states. This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
$\lambda^{\prime}=\bra{\mathbf{b}}\rho^{\prime}\ket{\mathbf{b}},$
The key update equation is
Appendix C Analysis of the Haar random input quantum states
In this section, we analyse the 2-qubit quantum states sampled from random_density_matrix of IBM qiskit’s quantum_info module. Here, we sample $100000$ Haar random quantum states and observe how the eigenvalues and the qubits’ conditional entropy change with the state’s entanglement.
In Fig. 8, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state. Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases. Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00. It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig. 1.
Sampling a random quantum state
If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module (in python) as follows
@@ Appendix C Analysis of the Haar random input quantum states
The state used in III.1, is sampled using $n=2$ and $s=27$

### v3

@@ V.2.1 Contribution of individual qubits for change in entanglement
We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Fig. 6. We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$. This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement. In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases. However, no such pattern is observed for the region of $\Delta_{q_{0}}^{q_{1}}<0$. When $\Delta_{c}>0.4$, the value of $\Delta_{q_{0}}^{q_{1}}$ monotonically decreases with increase in $\Delta_{c}$. Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
V.2.2 Problem with variational quantum state diagonalisation
@@
We present the main observations in Fig. 7, where we plot the four eigenvalues of $35$, $2$-qubit Haar random mixed quantum states with the correlation in conditional entropy among the qubits. For an individual quantum state, the RL-agent proposes $\mathbf{s}$ different structures of the admissible ansatzes. Then we get $\rho_{i}^{\prime}$ states where $i\in[0,\mathbf{s}]$. Afterwards, we calculate the conditional entropy of individual qubits for each state and take the median. [^1] Calculating the median over the mean is justified by the broad range of conditional entropy distribution, which makes the median a more appropriate measure of central tendency. Then, we calculate the correction in conditional entropy among the qubits using PCC.
@@
In Fig. 7, we observe that a mild anti-correlation between the qubits is required to find the two largest eigenvalues. As the magnitude of the largest eigenvalues decreases, the mild anti-correlation turns into a strong correlation between the qubits. This indicates that the qubits require mild to strong correlation to achieve the two largest eigenvalues. Meanwhile, we require a mild to strong anti-correlation between the qubits to find the smallest eigenvalues of the same states. Then, as the magnitude of the smallest eigenvalue increases, the strong anti-correlation turns into mild anti-correlation.
@@
Figure 7: On average over $35$ random quantum states the correlation in conditional entropy between the first and second qubits decreases as the magnitude of the largest and second largest eigenvalues increases. Meanwhile, the same correlation between the qubits increases with the magnitude of the third largest and the smallest eigenvalues. The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have both types of correlations simultaneously. Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues using the same ansatz that provides a good approximation of the largest eigenvalues. The correlation of conditional entropy between qubits is calculated using the Pearson correlation coefficient (PCC) and is the median over all configurations of RL-ansatz proposed by the agent. The $\bigstar$ signifies the eigenvalues corresponding to the same state.
@@
We recall that the correlation is calculated among the qubits based on the conditional entropy of individual qubits. As it is impossible to have two different kinds of correlation (such as mild correlation and mild anti-correlation simultaneously) between the qubits in the same ansatz structure, we can not find the largest and smallest eigenvalues utilising the same ansatz. For example, we mark the eigenvalues of a state with $\bigstar$, where we observe that to find the largest two eigenvalues, the RL-ansatz need to induce a mild correlation between the two qubits; meanwhile, to obtain the smallest eigenvalues, we require mild anti-correlation.
@@
Figure 8: The average correlation in conditional entropy among qubits is in the range of mild correlation for $35$ random quantum states for the first two largest eigenvalues whereas the conditional entropy among qubits requires a mild anti-correlation for the smallest eigenvalues. This observation, along with the observation in Fig. 7 certify that with a specific configuration of an admissible ansatz, we can not simultaneously find the largest and the smallest eigenvalues.
The average correlation in conditional entropy among qubits falls within the range of mild correlation for the first two largest eigenvalues across $35$ random quantum states. Conversely, the conditional entropy among qubits exhibits a mild anti-correlation for the smallest eigenvalues. This finding, together with the observation in Fig. 8, indicates that with a specific configuration of an admissible ansatz, it is interactable to simultaneously identify the largest and smallest eigenvalues. It should be noted that for the $35$ random quantum states, the average correlation in conditional entropy lies in the mild correlation zone, indicating that RL-VQSD can efficiently find the largest few eigenvalues. However, it fails to provide a good approximation for the smallest eigenvalues simultaneously.
VI Discussion
@@
We observe that the concurrence of the admissible ansatz ranges between an upper and lower bound, which are independent of the initialisation of the weights of the deep-Q network. The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement. The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase transition point at concurrence $0.322$. We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound. These observations provide insight regarding the relation between the entanglement of the initial state and the entanglement generated by the RL-VQSD admissible ansatzes. We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem. Specifically, we propose an entanglement enhancing block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$. Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
@@
Furthermore, we quantify the contribution of each qubit in the admissible ansatz using conditional quantum entropy. For the admissible RL-VQSD ansatz, we observe an equal contribution from each qubit for the full range of change in the entanglement of the state due to the ansatz. Focusing on the task of obtaining the eigenvalues of the starting state, we observe that a mild correlation between the qubits of the admissible ansatz is required to obtain the two largest eigenvalues, while mild anti-correlation between the qubits is needed to find the smallest eigenvalues of the same state. Noting that it is impossible to have both mild correlation and anti-correlation between the qubits of the same ansatz, we cannot simultaneously obtain both the largest and smallest eigenvalues from the RL-VQSD admissible ansatz. From the observation, most of the admissible ansatz has mild correlation among the qubits, which explains why it is easier for the VQSD algorithm to find the largest eigenvalues than the smallest ones.
@@
For future work, one could conduct a comprehensive study of the implication of introducing correlations such as entanglement Horodecki et al. (2009), nonlocality Brunner et al. (2014); Sadhu and Das (2023) or steering Uola et al. (2020); Sadhu and Das (2024)-based reward functions of the RL-agent in QAS. Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations. Also, we may expect this to provide an optimal quantum circuit that provides both the eigenvalues of a random quantum state with higher accuracy within a smaller time.
In this work, we present an intuitive explanation of the phase transition behavior of the bounds on the entanglement of the admissible ansatz. However, we point out that the phase transition point is problem-specific. It would be interesting to see if the phase transition behavior still holds for other classes of algorithms, a problem we leave for future research.
Furthermore, using a similar approach, presented in our paper one can rigorously investigate the RL-ansatz in the task of minimizing the entanglement entropy of a quantum many-body system to identify quantum phase transition An et al. (2021) and finding the ground state of chemical hamiltonian Patel et al. (2024); Ostaszewski et al. (2021).
@@ A.2 Conditional quantum entropy
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, consider a purification |$\ket{\psi}_{ABE}\in\mathscr{H}_{A}\otimes\mathscr{H}_{B}\otimes\mathscr{H}_{E}$. It then follows that:
@@ A.3 Correlation coefficient
A.4 Sampling a random quantum state
If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module as follows
@@ A.4 Sampling a random quantum state
The state used in V.1, is sampled using $n=2$ and $s=27$.